OpenStudy (anonymous):

If x^a = y^b = z^c and y^2 = xy then show that, 1/a + 1/c = 2/b

ganeshie8 (ganeshie8):

you mean y^2 = xz ?

OpenStudy (anonymous):

put x^a=y^b=z^c=t calculate x,y,z in terms of t and then put the values of x,y,z

OpenStudy (anonymous):

you followed or should i give you more clue.

OpenStudy (anonymous):

cud u xplain me in detail? @surjithayer

OpenStudy (anonymous):

\[x ^{a}=t,x=t ^{\frac{ 1 }{ a }}\] similarly y=? and z=?

OpenStudy (anonymous):

clear

OpenStudy (anonymous):

x^a = y^b = z^c = k x = k^(1/a) y = k^(1/b) z=k^(1/c) y^2 = x*z k^(2/b) = k^ (1/a + 1/b) --> 2/b = 1/a + 1/b

OpenStudy (anonymous):

soory *1/c

OpenStudy (anonymous):

\[x ^{a}=\left( t ^{\frac{ 1 }{a }} \right)^{a}=\left( t \right)^{\frac{ 1 }{ a }*a}=t\]